# Hunting New Animalcula

### v1

@@ 1 Introduction
While van Leeuvanhoek could reach the resolution down to roughly $10^{-6}$m, over the last 350 years this resolution could be improved by many orders of magnitude. On the way down to shortest distance scales scientists discovered nanouniverse ($10^{-9}$m), femtouniverse ($10^{-15}$m) relevant for nuclear particle physics and low energy elementary particle physics and finally attouniverse ($10^{-18}$m) that is the territory of contemporary high energy elementary particle physics.
@@
Another route is to study first the patterns of anomalies observed in the the data and compare them with the patterns of deviation from SM predictions in a given NP scenario. Such patterns, that expose suppressions and enhancements of various observables relative to SM predictions, can be considered as DNAs of the animalcula hunted by us. In particular the correlations between various enhancements and suppressions can rule out some NP scenarios already before any global fit is performed. This is the strategy proposed by Jennifer Girrbach-Noe and myself in 2013 Buras:2013ooa . It has been documented in several subsequent papers, in particular in my book Buras:2020xsm and recently in Buras:2024mhy with many colourful plots. Therefore I will not discuss it here.
@@
  • Tensions in the determinations of $|V_{cb}|$ and $|V_{ub}|$ from inclusive and exclusive tree-level decays Bordone:2021oof ; Finauri:2023kte ; FlavourLatticeAveragingGroupFLAG:2021npn . Using these results lowers the precision with which CKM parameters can be determined and consequently also lowers the precision of SM predictions for many observabies as illustrated soon. Therefore the inclusion of these determinations in the fit should be avoided until theorists agree what the values of $|V_{cb}|$ and $|V_{ub}|$ are.
@@ 2 BV-Strategy
In order to motivate this strategy in explicit terms, let us recall the values of $|V_{cb}|$ extracted from inclusive and exclusive tree-level semi-leptonic $b\to c$ decays Finauri:2023kte ; FlavourLatticeAveragingGroupFLAG:2021npn
@@
These uncertainties are clearly a disaster for those like me, my collaborators and other experts in NLO and NNLO calculations who spent decades to reduce theoretical uncertainties in basically all important rare $K$ and $B$ decays and quark mixing observables down to $(1-2)\%$ Buras:2011we . It is also a disaster for lattice QCD experts who for quark mixing observables and in particular meson weak decay constants achieved the accuracy at the level of a few percents FlavourLatticeAveragingGroupFLAG:2021npn .
@@ 5 CKM Parameters
$K^{+}\rightarrow\pi^{+}\nu\bar{\nu}$ $(8.60\pm 0.42)\cdot 10^{-11}$ $(13.0\pm 3.1)\cdot 10^{-11}$ NA62:2022hqi
@@
Table 1: Present most accurate SM estimates of the branching ratios for the the magnificant seven obtained using the BV-strategy. The experimental result on $K^{+}\rightarrow\pi^{+}\nu\bar{\nu}$ from NA62 includes the recent data presented in a seminar at CERN.
@@ 6 The Magnificant Seven
Until now we dealt only with the issue of the CKM parameters but the choice of the observables that are particularly theoretically clean is also very important. In my view the seven ones listed in Table 1 are particulary promising. We give there their SM values obtained using our strategy together with the present experimental results. Clearly other decays, in particular $B\to K\ell^{+}\ell^{-}$ and $B\to K^{*}\ell^{+}\ell^{-}$, that remain still central in the search for NP, are very important. Unfortunately, there are different views among theorists on the role of long distance QCD effects in these decays and the interpretation of the observed anomalies in these decays is difficult at present.
@@
In the case of B decays sufficient suppresion of NP contributions to $\Delta M_{s}$ and $\Delta M_{d}$ requires the presence of both left-handed and right-handed $Z^{\prime}$ couplings to quarks. The corresponding analysis for rare B decays has been performed in Buras:2024new finding again strong correlations between B-decay branching ratios in Table 1 and the ones for $B\to K\ell^{+}\ell^{-}$ and $B\to K^{*}\ell^{+}\ell^{-}$ allowing for the explanation of the anomalies in the latter ones despite the suppression of NP contributions to $\Delta F=2$ processes. In 331 models, in which only left-handed $Z^{\prime}$ couplings are involved, some amount of NP contributions at the level of $5\%$ has to be allowed to explain the anomalies in question Buras:2023ldz .
@@ 7 Summary and Shopping List
    Precise measurements of the branching ratios for the magnificant seven and of missing energy distributions for four decays among them. Presently on a forefront are the $B\to K^{(*)}\nu\bar{\nu}$ decays studied intensively by the Belle II experiment Belle-II:2023esi , giving some hints for NP contributions. For selected recent analyses of these data see Bause:2021cna ; He:2021yoz ; Bause:2022rrs ; Becirevic:2023aov ; Bause:2023mfe ; Allwicher:2023syp ; Dreiner:2023cms ; Altmannshofer:2023hkn ; Gabrielli:2024wys ; Hou:2024vyw ; He:2024iju ; Bolton:2024egx ; Marzocca:2024hua ; Allwicher:2024ncl . The same applies to $K^{+}\rightarrow\pi^{+}\nu\bar{\nu}$ for which a very interesting result has been provided very recently by the NA62 experiment listed in Table 1. Taking the present experimental result for $B_{s}\to\mu^{+}\mu^{-}$ allows us to determine the present $|V_{cb}|$-independent experimental values for the the ratios $R_{1}$ and $R_{7}$ in (8) and (14) that read
@@
    Clarification of anomalies in $B\to\pi K$ decays Buras:2003dj ; Buras:2004ub ; Fleischer:2018bld ; Berthiaume:2023kmp ; Datta:2024zrl ; Szabelski:2024cem .

### v2

@@ 1 Introduction
While van Leeuvanhoek could reach the resolution down to roughly $10^{-6}$m, over the last 350 years this resolution could be improved by many orders of magnitude. On the way down to shortest distance scales scientists discovered nanouniverse ($10^{-9}$m), femtouniverse ($10^{-15}$m) relevant for nuclear physics and low energy elementary particle physics and finally attouniverse ($10^{-18}$m) that is the territory of contemporary high energy elementary particle physics.
@@
Another route is to study first the patterns of anomalies observed in the data and compare them with the patterns of deviation from SM predictions in a given NP scenario. Such patterns, that expose suppressions and enhancements of various observables relative to SM predictions, can be considered as DNAs of the animalcula hunted by us. In particular the correlations between various enhancements and suppressions can rule out some NP scenarios already before any global fit is performed. This is the strategy proposed by Jennifer Girrbach-Noe and myself in 2013 Buras:2013ooa . It has been documented in several subsequent papers, in particular in my book Buras:2020xsm and recently in Buras:2024mhy with many colourful plots. Therefore I will not discuss it here.
@@
  • Tensions in the determinations of $|V_{cb}|$ and $|V_{ub}|$ from inclusive and exclusive tree-level decays Bordone:2021oof ; Finauri:2023kte ; FlavourLatticeAveragingGroupFLAG:2024oxs . Using these results lowers the precision with which CKM parameters can be determined and consequently also lowers the precision of SM predictions for many observabies as illustrated soon. Therefore the inclusion of these determinations in the fit should be avoided until theorists agree what the values of $|V_{cb}|$ and $|V_{ub}|$ are.
@@ 2 BV-Strategy
In order to motivate this strategy in explicit terms, let us recall the values of $|V_{cb}|$ extracted from inclusive and exclusive tree-level semi-leptonic $b\to c$ decays Finauri:2023kte ; FlavourLatticeAveragingGroupFLAG:2024oxs
@@
These uncertainties are clearly a disaster for those like me, my collaborators and other experts in NLO and NNLO calculations who spent decades to reduce theoretical uncertainties in basically all important rare $K$ and $B$ decays and quark mixing observables down to $(1-2)\%$ Buras:2011we . It is also a disaster for lattice QCD experts who for quark mixing observables and in particular meson weak decay constants achieved the accuracy at the level of a few percents FlavourLatticeAveragingGroupFLAG:2024oxs .
@@
It should be mentioned that to obtain these values we have used $\beta=22.2(7)^{\circ}$. A new average from HFLAV HFLAV:2022esi reads $\beta=22.6(4)^{\circ}$ fully compatible with the one used by us but more accurate. In view of other larger uncertainties it is better to wait for future updates of $\beta$ from Belle II before all these ratios will be updated.
@@ 5 CKM Parameters
$K^{+}\rightarrow\pi^{+}\nu\bar{\nu}$ $(8.60\pm 0.42)\cdot 10^{-11}$ $(13.0\pm 3.1)\cdot 10^{-11}$ NA62:2022hqi ; NA62:2024pjp
@@
Table 1: Present most accurate SM estimates of the branching ratios for the the magnificant seven obtained using the BV-strategy.
@@ 6 The Magnificant Seven
Until now we dealt only with the issue of the CKM parameters but the choice of the observables that are particularly theoretically clean is also very important. In my view the seven ones listed in Table 1 are particulary promising. We give there their SM values obtained using our strategy together with the present experimental results. Clearly other decays, in particular $B\to K\ell^{+}\ell^{-}$ and $B\to K^{*}\ell^{+}\ell^{-}$, that remain still central in the search for NP, are very important. Unfortunately, there are different views among theorists on the role of long distance QCD effects in these decays and the interpretation of the observed anomalies in these decays is difficult at present. It should be noted that in the case of $K_{S}\to\mu^{+}\mu^{-}$ also long distance contributions are present but fortunately as demonstated in Dery:2022yqc , generalizing the ideas of DAmbrosio:2017klp , the short distance contribution can be extracted and measured.
@@
In the case of B decays sufficient suppresion of NP contributions to $\Delta M_{s}$ and $\Delta M_{d}$ requires the presence of both left-handed and right-handed $Z^{\prime}$ couplings to quarks. The corresponding analysis for rare B decays has been performed in Buras:2024mnq finding again strong correlations between B-decay branching ratios in Table 1 and the ones for $B\to K\ell^{+}\ell^{-}$ and $B\to K^{*}\ell^{+}\ell^{-}$ allowing for the explanation of the anomalies in the latter ones despite the suppression of NP contributions to $\Delta F=2$ processes333 In 331 models, in which only left-handed $Z^{\prime}$ couplings are involved, some amount of NP contributions at the level of $5\%$ has to be allowed to explain the anomalies in question Buras:2023ldz .. The most interesting result of the analysis in Buras:2024mnq is the finding that in this scenario the suppression of $B\to K(K^{*})\mu^{+}\mu^{-}$ branching ratios observed by the LHCb implies uniquely enhancements of $B\to K(K^{*})\nu\bar{\nu}$ branching ratios as hinted by the most recent Belle II data.
@@ 7 Summary and Shopping List
    Precise measurements of the branching ratios for the magnificant seven and of missing energy distributions for four decays among them. Presently on a forefront are the $B\to K^{(*)}\nu\bar{\nu}$ decays studied intensively by the Belle II experiment Belle-II:2023esi , giving some hints for NP contributions. For selected recent analyses of these data see Buras:2024mnq ; Bause:2021cna ; He:2021yoz ; Bause:2022rrs ; Becirevic:2023aov ; Bause:2023mfe ; Allwicher:2023xba ; Dreiner:2023cms ; Altmannshofer:2023hkn ; Gabrielli:2024wys ; Hou:2024vyw ; He:2024iju ; Bolton:2024egx ; Marzocca:2024hua ; Allwicher:2024ncl . The same applies to $K^{+}\rightarrow\pi^{+}\nu\bar{\nu}$ for which a very interesting result has been provided very recently by the NA62 experiment listed in Table 1. Taking the present experimental result for $B_{s}\to\mu^{+}\mu^{-}$ allows us to determine the present $|V_{cb}|$-independent experimental values for the the ratios $R_{1}$ and $R_{7}$ in (8) and (14) that read
@@
    Clarification of anomalies in $B\to\pi K$ decays Buras:2003dj ; Buras:2004ub ; Fleischer:2018bld ; Berthiaume:2023kmp ; Datta:2024zrl ; Szabelski:2024cem ; Altmannshofer:2024kxb .
@@
    The above statement has been made in the first version of our paper. Recently, a NNLO calculation for $\hat{B}_{K}$ has been performed in Gorbahn:2024qpe which decreased the error in the RBC-UKQCD result but moved it upwards: $\hat{B}_{K}=0.7600(53)$. Thus the story of $\hat{B}_{K}$ is not finished yet. However, what is quite amazing is the fact that $\hat{B}_{K}$ is so close to its strict large $N$ limit: $\hat{B}_{K}=0.75$. Presently, to my knowledge, only DQCD can explain the dynamics behind this result. It is the strong cancellation of the negative pseudoscalar meson contributions to $\hat{B}_{K}$ by the positive vector meson contributions in the context of the so-called meson evolution within DQCD Buras:2014maa .